Your concave plate with a mass of 2 kg is spinning on top the table with a velocity of 1.5
m/s for 2 seconds then your brother came and spins it more creating a velocity of 4 m/s
for 6 seconds. What is the angular acceleration of your concave when you are spinning it
and the angular acceleration after your brother spins it?
1
Expert's answer
2020-03-30T07:53:35-0400
"a_1=\frac{v_1}{t_1}=\frac{1.5}{2}=0.75\frac{m}{s^2}"

"a_2=\frac{v_2-v_1}{t_2}=\frac{4-1.5}{6}=0.42\frac{m}{s^2}"